OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = 5/2x + 2. A. y = 2/5x + 1 B. y = -5/2x -27/2 C. y = -2/5x - 3 D. y = 5/2x + 23/2

OpenStudy (anonymous):

What is the slope of this line: y = 5/2x + 2?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

is it like the same thing we did last time?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh. i forgot how to do this then

OpenStudy (anonymous):

have you seen this equation before: y=mx+b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

m is the slope is If you look at the equation y = 5/2x + 2, can you see how it's the same format?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

So if you compare it to the format of y=mx+b, and if "m" represents the slope, can you find the slope of y = 5/2x + 2?

OpenStudy (anonymous):

so is the answer D?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

But you are close.

OpenStudy (anonymous):

The question asked for a perpendicular line to y = 5/2x + 2

OpenStudy (anonymous):

That means that it will be negative inverse

OpenStudy (anonymous):

so it will look like -2/5

OpenStudy (anonymous):

ohhhhh okay

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

no

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

oh thank you. Im really thankful of you helping me. :)

OpenStudy (anonymous):

yep!